OpenStudy (anonymous):

a family has two cars. The first car has a fuel efficiency of 20 miles per gallon of gas and the second has a fuel efficiency of 30 miles per gallon of gas. During one particular week, the two cars went a combined total of 1150 miles for a total gas consumption of 45 gallons. How many gallons were consumed by each of the two cars that week? Thank you for your help!

OpenStudy (anonymous):

I know this is an easy question but I am horrible at math. If i could have help by someone explaining what I need to multiple or divide it would help me answer this question. I appreciate any help.

OpenStudy (anonymous):

I just I dont understand how to add this together to come up with the correct answer. How do you corollate X+Y into know what the total gallons were used by each car. I just cant get my brain to understand how to get the answer.

OpenStudy (perl):

first it helps to keep in mind, there are 4 unknowns

OpenStudy (anonymous):

So what would the unknowns be? I seriously like havent done math in 13 plus years and am beyond frustrated right now and trying to re-learn and understand it. Im sorry for asking so many questions.

OpenStudy (perl):

\[\red ? \]

OpenStudy (perl):

you can ask all the questions you want

OpenStudy (perl):

the more questions the better (thats how i figure out your difficulty)

OpenStudy (perl):

do you agree with this equation (# of gallons used) * (# miles / gal) = # miles driven

OpenStudy (perl):

lets do a slightly easier approach

OpenStudy (anonymous):

Im sorry. I just i see the numbers on the page and in my head i think okay this will be easy, but instead all i see is jumbled together. I try to understand it but it just looks like jumbled and my brain cant just understand how everything gets put together.

OpenStudy (perl):

Let x = # of miles driven by car 1 1150 - x = # of miles driven by car 2 , since they add up to 1150 miles

OpenStudy (perl):

sometimes my first explanation does not make sense, so you ignore what i said earlier (do a mental reboot)

OpenStudy (anonymous):

I am really bad at algebra. I kinda understand basic fractions and know multiplication but my brain is very slow on the x/y stuff if that makes sense?

OpenStudy (perl):

yes

OpenStudy (perl):

if you know the number of miles that car 1 has driven, you can find the number of miles of car 2 (car 1 miles ) + (car 2 miles) = 1150

OpenStudy (anonymous):

okay I get that. they equal total 1150. So how do you figure out how you split that up. . Half of 1150 is 575. But car 1 only does 20mpg while car two does 30.

OpenStudy (perl):

we don't the exact amount each car got (not necessarily split in the middle)

OpenStudy (perl):

we don't know the exact amount each car got (not necessarily split in the middle)

OpenStudy (anonymous):

Than how am I supposed to answer this question they want me to answer?

OpenStudy (perl):

Let x = # of miles driven by car 1 1150 - x = # of miles driven by car 2 do you agree so far

OpenStudy (anonymous):

okay yes. I agree. But how do you figure out the number of miles driven by car 1? to come up with the numbers of car 2?

OpenStudy (anonymous):

And see now im lost again. like i cant decipher it. hah whats ridic, is all i need is to pass one damn math class for my degree, on the damn deans list and I cant pass simple math.

OpenStudy (perl):

sorry i made a typo (thats probably why it didn't make sense )

OpenStudy (anonymous):

haha I doubt your typo stopped it from making sense for me. I just. Cant get it.

OpenStudy (perl):

if you have a 100 mile trip, and your car gets 1/4 gallon per mile

OpenStudy (anonymous):

25?

OpenStudy (perl):

yes :)

OpenStudy (perl):

you have used up 25 gallons of gas

OpenStudy (anonymous):

Okay I think I kinda get it? I think what happens is my brain sees the X and then i get confused so I have to learn in my head to say gallons instead of X to understand it.

OpenStudy (anonymous):

I still though on my own just dont get how you added that up to get the outcome. 30 1st. 15 2nd.

OpenStudy (perl):

1st car DATA: rate = 20 mi/gal ; # of gallons consumed = x gal distance = 20x miles ------------------- 2nd car DATA: 30 mi/gal # of gallons consumed= 45-x gal distance = 30(45-x) miles

OpenStudy (anonymous):

So it would be 30 the first and 15 the second. ?

OpenStudy (perl):

nope, these are still unknowns

OpenStudy (perl):

we do know that distance car 1 + distance car 2 = 1150 miles

OpenStudy (perl):

so we can substitute

OpenStudy (anonymous):

I guess I just dont understand than. If they are unknowns how is there an answer?

OpenStudy (perl):

20x + 30 (45 - x ) = 1150

OpenStudy (anonymous):

yeah? i know the miles but how do i get the actual answer for this question? is there no answer?

OpenStudy (perl):

there is an answer

OpenStudy (perl):

sorry if i confused you. i was just setting up the equations.

OpenStudy (perl):

1st car DATA: rate = 20 mi/gal ; # of gallons consumed = x gal distance_car_1 = 20x miles ------------------- 2nd car DATA: 30 mi/gal # of gallons consumed= 45-x gal distance_car_2 = 30(45-x) miles distance_car_1 + distance_car_2 = 20x + 30(45-x) but we know total distance of the two cars is 1150 miles 1150 = 20x + 30(45-x)

OpenStudy (anonymous):

so what is the answer to it? how do these numbers create the answer?

OpenStudy (perl):

now you solve for x

OpenStudy (anonymous):

x? is the gallons?

OpenStudy (anonymous):

Would the answer be 25 for the first car and 20 for the second?

OpenStudy (perl):

yes x is the number of gallons consumed by car 1

OpenStudy (perl):

and x = 20